# Infinitely many pairs of non-isomorphic elliptic curves sharing the same BSD invariants

### v1

@@ 5. Infinitely many pairs of elliptic curves sharing the same BSD data
If $E$ is an elliptic curve over $\mathbb{Q}$, then there are only finite number of square-free integers $D$ such that $E_{D}$ has a torsion point of order $>2$.
@@
Since $\phi\circ\hat{\phi}=[p]$ and $\hat{\phi}\circ\phi=[p]$, we have $F(\phi)\circ F(\hat{\phi})=[p]$ and $F(\hat{\phi})\circ F(\phi)=[p]$. For a prime $l\neq p$ and for all $m$, since $\gcd(p,l)=1$, $F([p])=p:F(E_{1})[l^{\infty}]\to F(E_{2})[l^{\infty}]$ is bijective, it follows that $F(\phi):F(E_{1})[l^{\infty}]\to F(E_{2})[l^{\infty}]$ is an isomorphism. Since $F(E_{1})$ and $F(E_{2})$ are torsion groups, we have
@@
Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous via a degree–$2$ isogeny defined over $\mathbb{Q}$. Assume that
@@
By Theorem 4.1, the Kodaira symbols are the same. By Theorem 4.2, Tamagawa numbers at $l\geq 3$ are the same because $\Delta_{E_{1}}=\Delta_{E_{2}}$. For the period, $\Omega_{E_{1}}=\Omega_{E_{2}}$ follows from Proposition 4.3. Because $\mathrm{rank}(E_{i}/\mathbb{Q})=0$, $\mathrm{Reg}(E_{i}/\mathbb{Q})=1$ for $i\in\{1,2\}$.
@@
Examples of pairs $(E_{a},E_{b})$ of elliptic curves that satisfy conditions (1), (2), and (3) of Proposition 5.4 are $(38025.ck1,38025.ck2)$ (this is the same as the pair described in Proposition 6.1), $(38025.t1,38025.t2)$, $(38025.u1,38025.u2)$, $(38025.cl1,38025.cl2)$, $(207025.bf1,207025.bf2)$, $(207025.bu1,207025.bu2)$, and $(207025.cj1,207025.cj2)$. As explained in Remark 5.8, these are unique pairs of 2-isogenous elliptic curves that share the same minimal discriminants, up to twists.
@@
Suppose that $E_{1}(\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}$ and $E_{1}/\mathbb{Q}$ has good reduction at $2$. Suppose $(E_{1},E_{2})$ is a pair of a balanced isogenous elliptic curves with $E_{1}(\mathbb{Q})[2^{\infty}]\cong E_{2}(\mathbb{Q})[2^{\infty}]$, and $\Delta_{E_{1}}=\Delta_{E_{2}}$. Additionally, suppose $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}/\mathbb{Q})=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}/\mathbb{Q})$. Then there exist infinitely many square-free integers $D$ such that $\text{BSD}(E^{D}_{1}/\mathbb{Q})=\text{BSD}(E^{D}_{2}/\mathbb{Q})$ and Kodaira symbols at every prime and minimal discriminants are the same.
@@
For condition (1), For a square-free integer $D$, $E_{1}^{D}(\mathbb{Q})[2]\cong E_{2}^{D}(\mathbb{Q})[2]$. By Theorem 5.1, there are finitely many $D$ such that $E^{D}_{1}(\mathbb{Q})_{\text{tor}}$ or $E^{D}_{2}(\mathbb{Q})_{\text{tor}}$ acquire points of order greater than 4. Thus, for almost all square-free integers $D$, we have $E^{D}_{1}(\mathbb{Q})[2^{\infty}]\cong E^{D}_{2}(\mathbb{Q})[2^{\infty}]$. By Proposition 5.3 (2), $E^{D}_{1}(\mathbb{Q})\cong E^{D}_{2}(\mathbb{Q})$.
@@
For (3), let us take $(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$. Note that $E^{D}_{1}/\mathbb{Q},E^{D}_{2}/\mathbb{Q}$ is global minimal when the model of $E_{1}/\mathbb{Q},E_{2}/\mathbb{Q}$ is global minimal because $(D,\Delta_{E_{1}}\Delta_{E_{2}})=1$. Also, $E^{D}_{1}$ has good reduction at $2$. Since $\frac{\Delta_{E^{D}_{2}}}{\Delta_{E^{D}_{1}}}=\frac{\Delta_{E_{2}}}{\Delta_{E_{1}}}=1$, we have $\Delta_{E^{D}_{1}}=\Delta_{E^{D}_{2}}$.
@@
The pairs in Example 5.5 satisfy the hypothesis of Proposition 5.4. By Theorem 5.6, infinitely many pairs $(E^{D}_{a},E^{D}_{b})$ are the pairs we have sought. ∎
@@
In [1], pairs of isogenous elliptic curves over $\mathbb{Q}$ with the same minimal discriminants are called discriminant twins, and examples of these are classified in table 5 of [1]. Discriminant twins over $\mathbb{Q}$ are extremely rare. Recent work [1] proved that there is only one pair of discriminant twins among 2-isogenous curves up to twists: $(4225.h1,4225.h2)$. The curve 4225.h1 is a twist of our 38025.ck1 by $-195$. For distinct square–free integer $D_{1},D_{2}$, $E^{D_{1}},E^{D_{2}}$ do not have the same BSD invariants since the real period $\Omega_{E^{D_{1}}}=\frac{\Omega_{E}}{\sqrt{|D_{1}|}},\Omega_{E^{D_{2}}}=\frac{\Omega_{E}}{\sqrt{|D_{2}|}}$ never coincide.
@@ 6. Explicit infinite families of pairs
In this chapter, independently of Smith’s results, we give a concrete infinite family using classical 2-descent (see for example, [10], X, Remark 3.7 and Proposition 4.9) with a balanced isogeny case. The key point is that since this is a balanced isogeny case, we can simultaneously control $\Sha(E_{1}/\mathbb{Q})$ and $\Sha(E_{2}/\mathbb{Q})$ by using Remark 3.3.
@@ 6.1. Two descent for balanced isogenous pairs of elliptic curves
Let $E_{1}:y^{2}=x^{3}+25350x^{2}+2471625x$ and $E_{2}:y^{2}=x^{3}-50700x^{2}+632736000x$ (labeled as (38025.ck1, 38025.ck2) in LMFDB). Take a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{3}\right)=-1$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, and $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$. For such $D$, $\Sha(E_{1}^{D}/\mathbb{Q})[2]=0$ and $\Sha(E_{2}^{D}/\mathbb{Q})[2]=0$.
@@
$S\coloneqq\{v\mid v\text{ divides }\deg\phi\}\cup\{v\mid v\text{ divides }\Delta_{E_{1}}\Delta_{E_{2}}\}\cup\{\infty\}=\{2,3,5,13,D,\infty\}$
@@
$\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})\cong\left\{\bar{d}\in\mathbb{Q}(S,2)\;\middle|\;\begin{aligned} &\bar{C}_{d}(\mathbb{Q}_{v})\neq\emptyset,\forall v\in S,\text{where }\ \bar{C}_{d}\ \text{is the projective closure of}\\ &\text{affine curve}\ C_{d}:dy^{2}=d^{2}-50700dDx^{2}+632736000D^{2}x^{4}\ \text{in}\ \mathbb{P}^{3}\end{aligned}\right\}$
@@
$\bar{C}_{d}\,(d<0)$infty ±$\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{R}\ \text{since}\ d<0$
$\bar{C}_{d}\,(3\mid d)$ 3 ±$\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{Q}_{3}\ \text{since}\ v_{3}(\frac{632736000}{d})=2-n_{2}=1$
$\bar{C}_{2}$ 5 ±$\pm\sqrt{\frac{632736000}{2}}\notin\mathbb{Q}_{5}$
$\bar{C}_{13}$ 5 ±$\pm\sqrt{\frac{632736000}{13}}\notin\mathbb{Q}_{5}$
$\bar{C}_{D}$ 5 ±$\pm\sqrt{\frac{632736000}{D}}\notin\mathbb{Q}_{5}$
$\bar{C}_{5D}$ 13 ±$\pm\sqrt{\frac{632736000}{5D}}\notin\mathbb{Q}_{13}$
$\bar{C}_{2D}$ 13 ±$\pm\sqrt{\frac{632736000}{2D}}\notin\mathbb{Q}_{13}$
@@
$\bar{C}_{10}$ 2 ±$\pm\sqrt{\frac{632736000}{10}}\notin\mathbb{Q}_{2}$
When $d<0$, $\bar{C}_{d}\notin\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$. Thus, $n_{0}=0$. For $n_{2}\geq 1$, $v_{3}(dy^{2})=1+2v_{3}(y)$ is odd while $v_{3}(d^{2}-50700Ddx^{2}+632736000D^{2}x^{4})$ is even. Indeed, $v_{3}(d^{2})=2$, $v_{3}(50700Dd^{\prime}x^{2})=2+2v_{3}(x)$, $v_{3}(632736000D^{2}x^{4})=2+4v_{3}(x)$. When $v_{3}(x)$ is positive or negative, it is even. When $v_{3}(x)=0$, let $f(x)\coloneqq d^{2}-50700Ddx^{2}+632736000D^{2}x^{4}$ and $d=3d^{\prime}$. Now, $\frac{f(x)}{9}\in\mathbb{Z}_{3}$ and $\frac{f(x)}{9}=d^{\prime 2}-16900Dd^{\prime}x^{2}+70304000D^{2}x^{4}\equiv d^{\prime 2}+2Dd^{\prime}x^{2}+2D^{2}x^{2}\equiv 1+2Dd^{\prime}+2D^{2}\not\equiv 0\bmod 3$, so $v_{3}(f(x))=2$. Thus, $n_{2}$ should be 00. So, the candidate of $d\in\mathbb{Q}(S,2)$ such that $d\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$ is of the form $2^{n_{1}}5^{n_{3}}13^{n_{4}}D^{n_{5}}$. Thus, $d=2^{n_{1}}5^{n_{3}}13^{n_{4}}D^{n_{5}}$. The left candidates of $d$ is $d=2,5,13,65,130,D,5D,13D,65D,10,26,10D,26D,130D$.
To check these candidates, it is sufficient to prove $d=2,13,D,5D,2D,10D,10$ does not belong to $\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$. Indeed,
$d\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})\Leftrightarrow 65d\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$
holds since $65\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$(see (6.2)).
By the following Lemma 6.3, we have $\text{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})=\{1,65\}\cong\mathbb{Z}/2\mathbb{Z}.$
2-descent shows that $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{17}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})=0$. In Remark 6.4, let us take $D_{0}=17$ so that $17D\in{\mathbb{Q}_{v}^{\times}}^{2}$. This condition is equivalent to $\left(\frac{D}{3}\right)=-1$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, and $D>0$. Then, by Remark 6.4, since $\phi$ is balanced, we have $\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}$.
@@
[…]
@@
For $d=10D$, take $p(10D)=D$. $C_{10D}:y^{2}=10D-50700Dx^{2}+63273600Dx^{4}$. When $v_{D}(x)=0$, we claim $v_{D}(g(x))=0$ where $g(x)\coloneqq 10-50700x^{2}+63273600x^{4}$. The discriminant $\Delta$ of the polynomial $1063273600t^{2}-50700t+$ 10 is $\Delta=2^{4}\times 3^{2}\times 5^{3}\times 13^{3}$. Since $\left(\frac{\Delta}{D}\right)=\left(\frac{5\times 13}{D}\right)=\left(\frac{5}{D}\right)\left(\frac{13}{D}\right)=-1$, $v_{D}(g(x))=0$ because $g(x)\equiv 0\bmod D$ does not have solutions. Thus, $v_{D}(Dg(x))=1$ is odd. When $v_{D}(x)$ is positive or negative, since $v_{D}(f_{10D}(x))$ is odd, we have $C_{10D}(\mathbb{Q}_{D})=\emptyset$.
For $d=10$, take $p(10)=2$. Then $C_{10}:y^{2}=10-50700Dx^{2}+63273600D^{2}x^{4}$ and we have $v_{2}(y^{2})=2v_{2}(y),v_{2}(50700Dx^{2})=2+2v_{2}(x)$, and $v_{2}(63273600D^{2}x^{4})=7+4v_{2}(x)$. When $v_{2}(x)=-1$, by substituting $Z=2x$, $C_{10}:y^{2}=10-12675DZ^{2}+3954600D^{2}Z^{4}$ and $v_{2}(Z)=0$. Since $y^{2}\equiv 2-3\times 1\times 1\equiv 7\bmod 8$ does not hold. When $v_{2}(x)=-2$, by substituting $W=4x$, $C_{10}:4y^{2}=40-12675DW^{2}+988650D^{2}W^{4}$. Since $4y^{2}\equiv 0-3\times 1\times 1+2\equiv 7\bmod 8$, $v_{2}(x)$ is not $-2$. The left case is $v_{2}(x)\leq-2$ or $v_{2}(x)\geq 0$. Since $v_{2}(10-12675DZ^{2}+3954600D^{2}Z^{4})$ is odd, we have $C_{10}(\mathbb{Q}_{2})=\emptyset$. ∎

### v2

@@ 5. Infinitely many pairs of elliptic curves sharing the same BSD data
Let $E$ be an elliptic curve over $\mathbb{Q}$. There are at most finitely many square-free integers $D$ for which $E^{D}(\mathbb{Q})$ contains a torsion point of order >2>2.
@@
Since $\phi\circ\hat{\phi}=[p]$ and $\hat{\phi}\circ\phi=[p]$, we have $F(\phi)\circ F(\hat{\phi})=[p]$ and $F(\hat{\phi})\circ F(\phi)=[p]$. For a prime $\ell\neq p$ and for all $m$, since $\gcd(p,\ell)=1$, $F([p])=p:F(E_{1})[\ell^{\infty}]\to F(E_{1})[\ell^{\infty}]$ is bijective, it follows that $F(\phi):F(E_{1})[\ell^{\infty}]\to F(E_{2})[\ell^{\infty}]$ is an isomorphism. Since $F(E_{1})$ and $F(E_{2})$ are torsion groups, we have
@@
Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous via a degree $2$ isogeny defined over $\mathbb{Q}$. Assume that
@@
By Theorem 4.1, the Kodaira symbols are the same. By Theorem 4.2, Tamagawa numbers at $\ell\geq 3$ are the same because $\Delta_{E_{1}}=\Delta_{E_{2}}$. For the period, $\Omega_{E_{1}}=\Omega_{E_{2}}$ follows from Theorem 4.3. Because $\mathrm{rank}(E_{i}/\mathbb{Q})=0$, $\mathrm{Reg}(E_{i}/\mathbb{Q})=1$ for $i\in\{1,2\}$.
@@
An example of the pair $(E_{1},E_{2})$ that satisfies the conditions (1), (2), and (3) is $(y^{2}=x^{3}+25350x^{2}+2471625x,y^{2}=x^{3}-50700x^{2}+632736000x)$. The LMFDB labels of $E_{1}$ and $E_{2}$ are 38025.ck1 and 38025.ck2, respectively. Indeed, for $(1)$, $E_{1}(\mathbb{Q})[2^{\infty}]\cong E_{2}(\mathbb{Q})[2^{\infty}]$. For $(2)$, $\Sha(E_{1}/\mathbb{Q})[2]=\Sha(E_{2}/\mathbb{Q})[2]=0$ and thus $\Sha(E_{1}/\mathbb{Q})[2^{\infty}]\cong\Sha(E_{2}/\mathbb{Q})[2^{\infty}]=0$. For $(3)$, $\Delta_{E_{1}}=\Delta_{E_{2}}=3^{6}\cdot 5^{9}\cdot 13^{9}$. As explained in Remark 5.8, this is a unique pair of 2 isogenous elliptic curves that shares the same minimal discriminants, up to twists.
@@
Suppose $E_{i}/\mathbb{Q}\ (i\in\{1,2\})$ is an elliptic curve with $E_{i}(\mathbb{Q})[2]\simeq\mathbb{Z}/2\mathbb{Z}$. Suppose that $E_{1}$ and $E_{2}$ are isogenous via a degree $2$ balanced isogeny $\phi$. Suppose that $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}/\mathbb{Q})=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}/\mathbb{Q})$. Also, suppose that $E_{1}$ has good reduction at $2$ and $\Delta_{E_{1}}=\Delta_{E_{2}}$. Then, there exist infinitely many square-free integers $D$ such that $\text{BSD}(E^{D}_{1}/\mathbb{Q})=\text{BSD}(E^{D}_{2}/\mathbb{Q})$ and Kodaira symbols at every prime and minimal discriminants are the same.
@@
For condition (1), For a square-free integer $D$, $E_{1}^{D}(\mathbb{Q})[2]\cong E_{2}^{D}(\mathbb{Q})[2]$. By Theorem 5.1, there are finitely many $D$ such that $E^{D}_{1}(\mathbb{Q})_{\text{tor}}$ or $E^{D}_{2}(\mathbb{Q})_{\text{tor}}$ acquire points of order greater than 2. Thus, for almost all square-free integers $D$, we have $E^{D}_{1}(\mathbb{Q})[2^{\infty}]\cong E^{D}_{2}(\mathbb{Q})[2^{\infty}]$. By Proposition 5.3 (1), $E^{D}_{1}(\mathbb{Q})\cong E^{D}_{2}(\mathbb{Q})$.
@@
For (3), let us take $(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$. Note that $E^{D}_{1}/\mathbb{Q},E^{D}_{2}/\mathbb{Q}$ are global minimal when the model of $E_{1}/\mathbb{Q},E_{2}/\mathbb{Q}$ are global minimal because $(D,\Delta_{E_{1}}\Delta_{E_{2}})=1$. Also, $E^{D}_{1}$ has good reduction at $2$. Since $\frac{\Delta_{E^{D}_{2}}}{\Delta_{E^{D}_{1}}}=\frac{\Delta_{E_{2}}}{\Delta_{E_{1}}}=1$, we have $\Delta_{E^{D}_{1}}=\Delta_{E^{D}_{2}}$.
@@
The pairs in Example 5.5 satisfy the hypothesis of Proposition 5.4 and Theorem 5.6. By Theorem 5.6, infinitely many pairs $(E^{D}_{1},E^{D}_{2})$ are the pairs we have sought. ∎
@@
In [1], pairs of isogenous elliptic curves over $\mathbb{Q}$ with the same minimal discriminants are called discriminant twins, and examples of these are classified in table 5 of [1]. Discriminant twins over $\mathbb{Q}$ are extremely rare. Recent work [1] proved that there is only one pair of discriminant twins among 2-isogenous curves up to twists: $(4225.h1,4225.h2)$. The curve 4225.h1 is a twist of our 38025.ck1 by 195-195. For distinct square–free integer $D_{1},D_{2}$, $E^{D_{1}},E^{D_{2}}$ do not have the same BSD invariants since the real period $\Omega_{E^{D_{1}}}=\frac{\Omega_{E}}{\sqrt{|D_{1}|}},\Omega_{E^{D_{2}}}=\frac{\Omega_{E}}{\sqrt{|D_{2}|}}$ never coincide.
@@ 6. Explicit infinite families of pairs
In this chapter, independently of Smith’s results, we give a concrete infinite family using classical 2-descent (see for example, [11], X, Remark 3.7 and Proposition 4.9) with a balanced isogeny case.
@@ 6.1. Two descent for balanced isogenous pairs of elliptic curves
Let $E_{1}:y^{2}=x^{3}+25350x^{2}+2471625x$ and $E_{2}:y^{2}=x^{3}-50700x^{2}+632736000x$ (labeled as (38025.ck1, 38025.ck2) in LMFDB). Take a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, and $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$. For such $D$, $\Sha(E_{1}^{D}/\mathbb{Q})[2]=0$ and $\Sha(E_{2}^{D}/\mathbb{Q})[2]=0$.
@@
$S\coloneqq\{v\mid v\text{ divides }\deg\phi\}\cup\{v\mid v\text{ divides }\Delta_{E^{D}_{1}}\Delta_{E^{D}_{2}}\}\cup\{\infty\}=\{2,3,5,13,D,\infty\}$
@@
$\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})\cong\left\{\bar{d}\in\mathbb{Q}(S,2)\;\middle|\;\begin{aligned} &\bar{C}_{d}(\mathbb{Q}_{v})\neq\emptyset,\forall v\in S,\text{where }\ \bar{C}_{d}\ \text{is the projective closure of}\\ &\text{affine curve}\ C_{d}:dy^{2}=d^{2}-50700dDx^{2}+632736000D^{2}x^{4}\ \text{in}\ \mathbb{P}^{3}\end{aligned}\right\}$
@@
$\bar{C}_{d}\,(d<0)$ $\infty$ $\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{R}\ \text{since}\ d<0$
$\bar{C}_{d}\,(3\mid d)$ 3 $\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{Q}_{3}\ \text{since}\ v_{3}(\frac{632736000}{d})=2-n_{2}=1$
$\bar{C}_{2}$ 5 $\pm\sqrt{\frac{632736000}{2}}\notin\mathbb{Q}_{5}$
$\bar{C}_{13}$ 5 $\pm\sqrt{\frac{632736000}{13}}\notin\mathbb{Q}_{5}$
$\bar{C}_{D}$ 5 $\pm\sqrt{\frac{632736000}{D}}\notin\mathbb{Q}_{5}$
$\bar{C}_{5D}$ 13 $\pm\sqrt{\frac{632736000}{5D}}\notin\mathbb{Q}_{13}$
$\bar{C}_{2D}$ 13 $\pm\sqrt{\frac{632736000}{2D}}\notin\mathbb{Q}_{13}$
@@
$\bar{C}_{10}$ 2 $\pm\sqrt{\frac{632736000}{10}}\notin\mathbb{Q}_{2}$
When $d<0$, $\bar{C}_{d}\notin\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$. Thus, $n_{0}=0$. For $n_{2}\geq 1$, $v_{3}(dy^{2})=1+2v_{3}(y)$ is odd while $v_{3}(d^{2}-50700Ddx^{2}+632736000D^{2}x^{4})$ is even. Indeed, $v_{3}(d^{2})=2$, $v_{3}(50700Dd^{\prime}x^{2})=2+2v_{3}(x)$, $v_{3}(632736000D^{2}x^{4})=2+4v_{3}(x)$. When $v_{3}(x)$ is positive or negative, it is even. When $v_{3}(x)=0$, let $f(x)\coloneqq d^{2}-50700Ddx^{2}+632736000D^{2}x^{4}$ and $d=3d^{\prime}$. Now, $\frac{f(x)}{9}\in\mathbb{Z}_{3}$ and $\frac{f(x)}{9}=d^{\prime 2}-16900Dd^{\prime}x^{2}+70304000D^{2}x^{4}\equiv d^{\prime 2}+2Dd^{\prime}x^{2}+2D^{2}x^{2}\equiv 1+2Dd^{\prime}+2D^{2}\not\equiv 0\bmod 3$, so $v_{3}(f(x))=2$. Thus, $n_{2}$ should be 0. So, the candidate of $d\in\mathbb{Q}(S,2)$ such that $d\in\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$ is of the form $2^{n_{1}}5^{n_{3}}13^{n_{4}}D^{n_{5}}$. Thus, $d=2^{n_{1}}5^{n_{3}}13^{n_{4}}D^{n_{5}}$. The left candidates of $d$ is d=2,5,13,65,130,D,5D,13D,65D,10,26,10D,26D,130Dd=2,5,13,65,130,D,5D,13D,65D,10,26,10D,26D,130D.
To check these candidates, it is sufficient to prove $d=2,13,D,5D,2D,10D,10$ does not belong to $\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$. Indeed,
$d\in\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})\Leftrightarrow 65d\in\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$
holds since $65\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$(see (6.2)). By the following Lemma 6.3, we have $\text{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})=\{1,65\}\cong\mathbb{Z}/2\mathbb{Z}.$
By Remark 3.3, $\tau^{D}=\prod_{p}\frac{1}{2}\#W_{p}(\phi,D)$. For any prime $p\neq 2$, by Lemma 2.1 (ii) of [14], we have $\frac{1}{2}\#W_{p}(\phi,D)=\frac{c_{E_{1}/\mathbb{Q}_{p}}}{c_{E_{2}/\mathbb{Q}_{p}}}=1$ where the last equality holds by Theorem 4.2. For $p=2$, $\frac{1}{2}{\#W_{2}(\phi,D)}=1$ since $E^{D}/\mathbb{Q}$ has good reduction at $2$. At the infinite place $p=\infty$, by Lemma 2.1 (iii), we have $\frac{1}{2}\#W_{\infty}(\phi,D)=1$ since $2471625D^{2}>0$. Thus, we have $\text{dim}_{\mathbb{F}_{2}}\text{Sel}^{\phi}(E^{D}_{1}/\mathbb{Q})=\text{dim}_{\mathbb{F}_{2}}\text{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})$ since $\tau^{D}=1$.
@@
For $d=5D$, take $p(5D)=13$. $C_{5D}:y^{2}=5D-50\,700D\,x^{2}+126\,547\,200D\,x^{4}$. When $v_{13}(x)\geq 0$, $y^{2}\equiv 5D\bmod 13$. This contradicts that $\left(\frac{5D}{13}\right)=-\left(\frac{D}{13}\right)=-1$. When $v_{13}(x)<0$, since $v_{13}(f_{5D}(x))=\text{min}\{0,2+2v_{13}(x),3+4v_{13}(x)\}=3+4v_{13}(x)$ is odd, we have $C_{5D}(\mathbb{Q}_{13})=\emptyset$.
@@
For $d=10D$, take $p(10D)=D$. $C_{10D}:y^{2}=10D-50700Dx^{2}+63273600Dx^{4}$. When $v_{D}(x)=0$, we claim $v_{D}(g(x))=0$ where $g(x)\coloneqq 10-50700x^{2}+63273600x^{4}$. The discriminant $\Delta$ of the polynomial $63273600t^{2}-50700t+10$ is $\Delta=2^{4}\times 3^{2}\times 5^{3}\times 13^{3}$. Since $\left(\frac{\Delta}{D}\right)=\left(\frac{5\times 13}{D}\right)=\left(\frac{5}{D}\right)\left(\frac{13}{D}\right)=-1$, $v_{D}(g(x))=0$ because $g(x)\equiv 0\bmod D$ does not have solutions. Thus, $v_{D}(Dg(x))=1$ is odd. When $v_{D}(x)$ is positive or negative, since $v_{D}(f_{10D}(x))$ is odd, we have $C_{10D}(\mathbb{Q}_{D})=\emptyset$.
For $d=10$, take $p(10)=2$. Then $C_{10}:y^{2}=10-50700Dx^{2}+63273600D^{2}x^{4}$ and we have $v_{2}(y^{2})=2v_{2}(y),v_{2}(50700Dx^{2})=2+2v_{2}(x)$, and $v_{2}(63273600D^{2}x^{4})=7+4v_{2}(x)$. When $v_{2}(x)=-1$, by substituting $Z=2x$, $C_{10}:y^{2}=10-12675DZ^{2}+3954600D^{2}Z^{4}$ and $v_{2}(Z)=0$. Since $y^{2}\equiv 2-3\times 1\times 1\equiv 7\bmod 8$ does not hold. When $v_{2}(x)=-2$, by substituting $W=4x$, $C_{10}:4y^{2}=40-12675DW^{2}+988650D^{2}W^{4}$. Since $4y^{2}\equiv 0-3\times 1\times 1+2\equiv 7\bmod 8$, $v_{2}(x)$ is not 2-2. The left case is $v_{2}(x)\leq-2$ or $v_{2}(x)\geq 0$. Since $v_{2}(10-12675DZ^{2}+3954600D^{2}Z^{4})$ is odd, we have $C_{10}(\mathbb{Q}_{2})=\emptyset$. ∎